State Newton s law of cooling. If the temperature of the air is 30^{0}C and thetemperature of the body drops from 80°C to 60^{0}C in 10 minutes. What will beits temperature after 20 minutes.


1
Expert's answer
2021-08-28T06:15:59-0400

Newton's law is a physical law that states that the rate at which a substance is cooling is proportional to the difference in temperature between the object and the object's surroundings.


"T_a=30°C\\
 T_b=80°C \\
T_f=60°C\\
t=10min\\
(60–30)=(80–60) exp^{-10k}\\( exp^{-10k}=3/2)"


Let T be the temperature of the body 20 minutes thereafter.

"(T-30)=(60-T)exp^{-20k}\\ [exp^{-10k}]²=\dfrac{(T-30)}{(60-T)}\\ \dfrac{(T-30)}{(60-T)}=9/4 \\
T-30 = 9/4(60-T)\\
T-30 = 135 -9/4T\\
T+9/4T = 135+30\\
T× 13/4 = 165\\
T = 165 × 4/13
\\ 
T≈50.77°C."